OpenStudy (goformit100):

“The product of two consecutive positive integers is divisible by 2”. Is this statement true or false? Give reasons.

OpenStudy (goformit100):

@genius12

OpenStudy (anonymous):

2 consecutive positive integers means one integer must be even and the other must be odd all even numbers are divisible by 2 when you multiply the odd and even together, you can still divide out the 2 from the even number

OpenStudy (anonymous):

False. 1*3 = 3 <- not even lol.

OpenStudy (anonymous):

consecutive?

OpenStudy (goformit100):

Ok..

mathslover (mathslover):

True. There will be an even and odd, so of course their product will be divisible by 2.

OpenStudy (anonymous):

Yes ok guys I think @goformit100 gets the idea here. we don't all need to crowd around the same question. move on to other questions.

mathslover (mathslover):

Goformit, is it clear to you?

mathslover (mathslover):

You can take cases here. Case 1 : a is even. a = 2q (using euclid's division algorithm, q is positive integer) \(\bf{\cfrac{a(a+1)}{2} = \cfrac{2q(a+1)}{2} = q(a+1)}\) , which is a positive integer. So, a(a+1) is divisible by 2. Case 2 : a + 1 is even \(\bf{\cfrac{a(a+1)}{2} = \cfrac{2q(a)}{2} = aq}\) which is also an integer So, a(a+1) is divisible by 2.

OpenStudy (goformit100):

Thank you Users :)

mathslover (mathslover):

You are welcome :)